### v1

# Multi-Class Stackelberg Games for the Co-Design of Networked Systems

Abstract
We investigate a co-design problem, encompassing simultaneous design of system infrastructure and control, through a game-theoretical framework. To this end, we propose the co-design problem as a two-layer hierarchical strategic interaction. At the upper layer, a leader (or multiple leaders) determines system design parameters, while at the lower layer, a follower (or multiple followers) optimizes the control strategy. To capture this hierarchy, we propose four novel classes of Stackelberg games that integrate diverse strategic behaviors, including combinations of cooperative and non-cooperative interactions across two different layers. Notably, the leaders’ interactions are represented using a normal-form game, whereas the followers’ interactions are modeled by different games (dynamic games in discrete time). These distinct game structures result in a Stackelberg game that accommodates different game types per layer, and/or supports heterogeneous strategic behaviors involving cooperation and non-cooperation simultaneously. Learning algorithms using the best-response dynamics are used to solve the game problems when considering a discrete strategic space for the leaders. The efficacy of the proposed approach is demonstrated through an application to the co-design of the Barcelona drinking water network.
Index Terms:
Networked systems, Stackelberg games, dynamic games, normal-form games, optimal control, learning
I Introduction
Game theory allows the study and modeling of strategic interactions, and it has been extensively used in the solution of engineering problems, e.g., for power systems, robot coordination, water distribution systems, etc. Several classes of games have been studied and reported in the literature, capturing different situations, such as cooperation, competition, altruism, and co-opetition. One of the most well-known game-theoretical solution concepts is the Nash equilibrium, which describes a situation in which there are no incentives for a single decision-maker to modify his/her selection in an unilateral way. This solution concept takes place in the non-cooperative games [1]. However, more possibilities have been explored in the game theory community, e.g., Stackelberg games to describe major and minor decision-makers or hierarchical interactions [2], [3, Chapter 7], coalitional games to understand the power index and/or level of influence within a strategic interaction [4, 5], cooperative games describing common interest payoffs, Berge games to describe altruism [6], co-opetition to describe simultaneously levels of cooperation and competition [7], [3, Chapter 5], among others. Indeed, it has been extensively shown that game theory is a suitable and powerful theoretical tool to model non-centralized control strategies.
Modern networked systems are inherently complex and large-scale, comprising multiple interconnected subsystems that must coordinate to achieve overall system efficiency and reliability [8, 9]. Examples of such systems include water distribution networks [10, 11], power grids [12], supply chains [13], and transportation systems [14]. The design and control of such systems pose significant challenges, particularly in industrial applications where system infrastructure is typically designed based on conservative estimates, and advanced control strategies are developed afterward. This sequential approach often results in suboptimal performance, as the control strategy is constrained by a fixed system design. To address this issue, we focus on co-design problem, an approach that simultaneously optimizes both system infrastructure and the associated control strategy [15]. A solution of the co-design problem could achieve the global optimum regarding to the overall objectives.
In this work, we focus on a co-design problem, which is suitably modeled by considering the combination of game-theoretical solution concepts at different stages. From a global view, we consider two layers, which can be interpreted as the designer of the system and the entity in charge of its operation. Yet, we can study strategic interaction possibilities within each one of these layers. We pursue the study of how designer decisions can be modeled when multiple entities are making decisions at this level, or we might consider that this layer is composed of a unique entity performing the design and solving an optimization problem. Regarding the operator of the system, we might encounter multiple options for strategic behavior that mainly span centralized and non-centralized approaches.
The main contribution of this paper is to propose a two-layer hierarchical game or a Stackelberg-like game problem that combines at each layer different classes of games. For the leaders’ layer, we have a normal form non-cooperative game with a finite number of strategies and whose utility functions depend on both the leaders’ and followers’ strategic selection. Hence, at this leaders’ layer, we may also consider cooperation, in which we perform optimization for the design parameters of the system. Regarding the followers’ layer, we introduce a dynamic game, more precisely, a non-cooperative difference game problem. Similarly, we may also consider a cooperative game at the followers’ layer, which corresponds to a standard optimal control problem [16]. We highlight the fact that the interaction between the layers in the hierarchical scheme is different from others reported in the literature. The hierarchical scheme coupling is mainly given by the fact that the leader defines the feasible set of strategies for the follower. We show that the proposed Stackelberg game, which combines leader normal form games with discrete strategic sets and follower difference games (dynamic games), can be used for solving the co-design control and system problem. The four Stackelberg game classes that we study in this paper are:
  • Non-cooperative leaders and followers,
  • Cooperative leaders and followers,
  • Non-cooperative leaders and cooperative followers,
  • Cooperative leaders and non-cooperative followers,
which will be formally introduced and explained later on throughout the paper.
The remainder of this paper is organized as follows. Section II presents the preliminaries comprising the game settings for both the leader and follower decision-makers, game-theoretical solution concepts, and the price of anarchy in this co-design problem context. Section III introduces the proposed four Stackelberg game classes together with their corresponding Stackelberg equilibria. Section IV presents the networked system application we use to illustrate the contributions of this paper. In Section V, the results are presented and discussions are developed to compare the different Stackelberg game classes. Finally, concluding remarks and future directions are summarized in Section VI.
II Problem Statement
This paper considers a networked engineering system and aims to analyze two main problems using game-theoretical tools. On one hand, we analyze the design of the networked system led by a leader’s layer. At this design stage, some key parameters for the system are determined. On the other hand, we analyze the control design to operate the system led by a followers’ layer. At this design stage, optimal control policies are defined to operate the system subject to the established designed parameters by the leaders. There are two classes of decision-makers, i.e., leaders and followers. The set of $L\in\mathbb{N}$ leaders is given by :=$\mathcal{L}:=\{1,\dots,L\}$, and the set of $M\in\mathbb{N}$ followers is given by :=$\mathcal{M}:=\{1,\dots,M\}$. The leader decision-makers set takes care of the system design and parameter setting, whereas the second decision-makers set is in charge of designing the optimal control to operate the system, adhering to the imposed rules of the leader’s set. The leaders interact with each other using a static normal form game, and their decisions are coupled with the followers’ interactions. The followers interact with each other by means of a difference game whose settings are leader-strategic-dependent.
In the following, we explain the settings for each game layer, i.e., for the leaders and followers, and then we formally introduce the corresponding game problems together with the different game-theoretical solution concepts we are interested in. Let us start with the followers’ strategic interaction settings.
II-A Followers Strategic Interaction and Settings
The followers interact on a difference game problem within a discrete-time interval $[0..T]:=[0,T]\cap\mathbb{Z}_{+}$, with $T\in\mathbb{N}$, involving a dynamical system given as follows:
$x_{k+1}$ $=f(x_{k},\{u_{i,k}\}_{i\in\mathcal{M}}),~{}\forall~{}k\in[0..(T-1)],$ (1a)
$x_{0}$ $\in\mathbb{X}(\cdot)~{}\text{given},$ (1b)
[…]
$\mathcal{U}_{i}:=\{\mathbf{u}_{i}:=u_{i,0:T-1}:u_{i,k}\in\mathbb{U}_{i},~{}\forall~{}k\in[0..(T-1)]\}$
denote the set of admissible controls of the $i$-th decision-maker. Also, we denote $\mathbf{u}_{-i}$ as the strategic selection sequences along the time horizon for all the decision-makers different from $i$, i.e.,
$\mathbf{u}_{-i}$ $:=(\mathbf{u}_{1},\dots,\mathbf{u}_{i-1},\mathbf{u}_{i+1},\dots,\mathbf{u}_{M})\in\mathcal{U}=\prod_{j\in\mathcal{M}\setminus\{i\}}\mathcal{U}_{j},$ (2)
where $\mathbf{u}:=(\mathbf{u}_{1},\dots,\mathbf{u}_{M})$ is the strategic profile or the joint strategic actions for the $M$ follower decision-makers.
II-B Followers Non-Cooperative Behavior
Each decision-maker seeks to minimize its own cost functional given by $V_{i}(x_{0},\mathbf{u}_{i},\mathbf{u}_{-i})$, and the difference game problem $\mathscr{P}^{\mathrm{NC}}_{\mathrm{follower}}$ is expressed as follows:
$\mathscr{P}^{\mathrm{NC}}_{\mathrm{follower}}:~{}\forall~{}i\in\mathcal{M},$ (3)
[…]
where each cost $V_{i}$ is assumed to be continuous, convex and coercive in $\mathbf{u}_{i}$, for all $i\in\mathcal{M}$, such that the optimization problem is well-defined. Given that the follower layer game problem depends on the decisions made at the leader level, we make the following assumption.
Assumption 1
The optimization problem (3) is feasible for every leader strategic profile $\mathbf{a}\in\mathcal{A}$. $\square$
Once the follower’s game problem is defined, we introduce the equilibrium concept by using the best-response strategies. First, we present the best way a follower can react to a design in Definition 1.
Definition 1 (Best-response strategies among followers)
A feasible control strategy $\mathbf{u}_{i}\in\mathcal{U}_{i}$ is a best response strategy if it solves the Problem in (3), given the strategic selection by other decision-makers $\mathcal{M}\setminus\{i\}$ given by $\mathbf{u}_{-i}$ and the strategic selection of the leader(s). The set of best-response strategies for the $i$-th decision-maker is denoted by $\mathbb{BR}^{\mathcal{M}}_{i}(\mathbf{u}_{-i},\{\text{leader strategies}\})$. $\square$
Then, we use the best-response strategy concept to introduce the game-theoretical equilibrium concept in Definition 2 below.
Definition 2 (Nash equilibrium among followers)
A strategic profile $\mathbf{u}^{*}=(\mathbf{u}^{*}_{1},\dots,\mathbf{u}^{*}_{M})\in\mathcal{U}$ is a Nash equilibrium if all the strategies are best-response strategies against each other given the strategic selection of the leader(s), i.e., $\mathbf{u}^{*}_{i}\in\mathbb{BR}^{\mathcal{M}}_{i}(\mathbf{u}^{*}_{-i},\{\text{leader strategies}\})$, for all $i\in\mathcal{M}$. The set of Nash equilibria for the followers strategic interaction is denoted by $\mathbb{NE}_{\mathrm{follower}}$. $\square$
Later on, in Section , we develop some algorithms to find the Nash equilibria. Under these algorithms, one might arbitrarily converge to a point close to the Nash equilibrium. Thus, Definition 3 below introduces the concept of $\varepsilon$ equilibrium, in which we can finish the solution seeking.
Definition 3 ($\varepsilon$-Nash equilibrium among followers)
A strategic profile $\mathbf{u}^{*}=(\mathbf{u}^{*}_{1},\dots,\mathbf{u}^{*}_{M})\in\mathcal{U}$ is a $\varepsilon$-Nash equilibrium if
$V_{i}(x_{0},\mathbf{u}_{i}^{*},\mathbf{u}_{-i}^{*})\leq V_{i}(x_{0},\mathbf{u}_{i},\mathbf{u}_{-i}^{*})+\varepsilon,$ (4)
for all $\mathbf{u}_{i}\in\mathcal{U}_{i}$, and $i\in\mathcal{M}$. The set of $\varepsilon$-Nash equilibria is denoted by $\varepsilon$-$\mathbb{NE}_{\mathrm{follower}}$. $\square$
Definition 4 (Best-response of non-cooperative followers strategies against leaders’ strategies)
A feasible strategic follower profile $\mathbf{u}^{*}\in\mathcal{U}$ is a non-cooperative best response against the strategic selection of the leader strategy if $\mathbf{u}^{*}$ is a Nash equilibrium. We denote by $\mathbb{BR}^{\mathrm{NC}}_{\mathcal{M}}(\{\text{leader strategies}\})$ the set of non-cooperative best response strategies given the strategic selection by the leader, i.e., $\mathbf{u}^{*}\in\mathbb{BR}^{\mathrm{NC}}_{\mathcal{M}}(\{\text{leader strategies}\})$. $\square$
We also would like to study the case in which the followers interact in a cooperative framework. This strategic behavior is formally presented in the next subsection.
II-C Followers Cooperative Behavior
The followers can interact in a cooperative game by jointly optimizing a common cost functional. The cooperative game problem, which corresponds to a standard optimal control problem, is formulated as follows:
$\mathscr{P}^{\mathrm{C}}_{\mathrm{follower}}:$ (5)
[…]
For simplicity in the notation, let us consider
$V(x_{0},\mathbf{u})=\sum_{i\in\mathcal{M}}V_{i}(x_{0},\mathbf{u}_{i},\mathbf{u}_{-i}).$
Similar to the Assumption 1 stated in the non-cooperative game problem settings, we state the following assumption for the cooperative case.
Assumption 2
The optimization problem (5) is feasible for every leader strategic profile $\mathbf{a}\in\mathcal{A}$. $\square$
Next, we introduce the best action that a follower can take, in a cooperative way, against the decisions made by the leader, in Definition 5. Note that this is different from the Definition 1.
Definition 5 (Best-response of cooperative followers strategies against leaders’ strategies)
A feasible strategic follower profile $\tilde{\mathbf{u}}^{*}\in\mathcal{U}$ is a cooperative best response against the strategic selection of the leader strategy if it solves the Problem in (5), i.e.,
$\tilde{\mathbf{u}}^{*}(\{\text{leader strategies}\}$ $)\in\arg\min_{\mathbf{u}\in\mathcal{U}}\sum_{i\in\mathcal{M}}V_{i}(x_{0},\mathbf{u}_{i},\mathbf{u}_{-i}),$
s. t. {(1),$\ x_{k}\in\mathbb{X}(\{\text{leader strategies}\}),\\ \text{for all}~{}k\in[0..T],\end{cases}$ (6)
We denote $\mathbb{BR}^{\mathrm{C}}_{\mathcal{M}}(\{\text{leader strategies}\})$ as the set of cooperative best response strategies given the strategic selection by the leader. $\square$
Once the strategic interactions for the followers are introduced, we proceed to formally present the strategic interactions corresponding to the leaders.
II-D Leaders Strategic Interaction and Settings
The leaders action set is $\{A_{1},\dots,A_{L}\}$ and each set contains $N\in\mathbb{N}$ actions. The $j-$th leader chooses an action $a_{j}\in A_{j}$, and a strategic profile is given by the joint selection of strategies along all the players, i.e.,
$\mathbf{a}=(a_{1},\dots,a_{L})\in\mathcal{A}=\prod_{j\in\mathcal{L}}A_{j}.$
Each leader decision-maker $j\in\mathcal{L}$ has an associated cost function, denoted by $J_{j}(\mathbf{a},\cdot)=J_{j}(a_{j},a_{-j},\cdot)$, for all $j\in\mathcal{L}$, to be minimized. Note that the cost functional of the leader decision-makers depends on the whole leader strategic profile and other terms coming from the decisions made by the followers, i.e., $J_{j}(a_{j},a_{-j},\{\text{follower strategies}\})$.
II-E Leaders Non-Cooperative Behavior
To achieve a non-cooperative behavior, each leader makes decisions by solving the following optimization problem:
$\mathscr{P}^{\mathrm{NC}}_{\mathrm{leader}}:~{}\forall~{}j\in\mathcal{L},$
$\min_{a_{j}\in A_{j}}J_{j}(a_{j},a_{-j},\{\text{follower strategies}\}),$ (7)
where the cost functional $J_{j}$ is assumed to be continuous, convex, and coercive in the strategies $a_{j}$, for all $j\in\mathcal{L}$, such that the optimization problem is well defined. We next introduce the best decision made by a leader depends on the decisions made by all the other leaders.
Definition 6 (Best-response strategies among leaders)
A strategy $a_{j}\in A_{j}$ is a best response strategy if for a given $a_{-j}$ that is the optimal strategy that minimizes the cost of the $j-$th leader, i.e., if it solves the problem in (II-E). The set of best response strategies for the $j-$th leader is denoted by $\mathbb{BR}^{\mathcal{L}}_{j}(a_{-j},\{\text{follower strategies}\})$. $\square$
We introduce the game-theory equilibrium concept in Definition 7 by using the best-response strategy in Definition 6.
Definition 7 (Nash equilibrium among leaders)
A feasible strategic profile $\mathbf{a}^{*}:=(a_{1}^{*},\dots,a_{L}^{*})\in\mathcal{A}$ is a Nash equilibrium if no leader has incentives to unilaterally change its strategy,
[…]
$\leq J_{j}(a_{j},a^{*}_{-j},\{\text{follower strategies}\}),$ (8)
for all $a_{j}\in A_{j}$, and $j\in\mathcal{L}$. Alternatively, the strategic profile $\mathbf{a}^{*}:=(a_{1}^{*},\dots,a_{L}^{*})$ is a Nash equilibrium if all strategies are best-response strategies against each other, i.e., $a_{j}^{*}\in\mathbb{BR}^{\mathcal{L}}_{j}(a_{-j},\{\text{follower strategies}\})$, for all $j\in\mathcal{L}$. The set of Nash equilibria for the leaders strategic interaction is denoted by $\mathbb{NE}_{\mathrm{leader}}$. $\square$
There are two important aspects to mention related to the game settings for the leaders’ problem. On one hand, note that the strategic set is considered to the discrete, i.e., there is a finite number of possible decisions that the leaders can make. On the other hand, we are only evaluating the pure strategies for the leaders. Therefore, we need to set the following assumption.
Assumption 3
The normal form game problem presented in (II-E) admits a Nash equilibrium in pure strategies introduced in Definition 7. $\square$
Note that the relaxation of Assumption 3 is challenging as it implies obtaining an expression of the optimal solution for the followers in terms of the leader’s decision. The computation of this optimal representation, for the followers in terms of the leaders, is not that involved when the leaders’ actions affect either the dynamical system or the cost for the followers. In contrast, in this problem setting, the leaders’ decisions affect the feasible set of the system states.
In Section III, we will present how to approximate this solution by using an algorithm, in which we use an $\epsilon$ equilibrium condition that is introduced next in Definition 8.
Definition 8 ($\varepsilon$-Nash equilibrium among leaders)
A feasible strategic profile $\mathbf{a}^{*}:=(a_{1}^{*},\dots,a_{L}^{*})\in\mathcal{A}$ is an $\varepsilon$-Nash equilibrium if
[…]
$\leq J_{j}(a_{j},a^{*}_{-j},\{\text{follower strategies}\})+\varepsilon,$ (9)
for all $a_{j}\in A_{j}$, and $j\in\mathcal{L}$. The set of $\varepsilon$-Nash equilibria is denoted by $\varepsilon$-$\mathbb{NE}_{\mathrm{leader}}$. $\square$
To complete all the possible combinations in the interactions among leaders and followers, we also consider the case in which the leaders behave in a cooperative way as introduced in the following subsection.
II-F Leaders Cooperative Behavior
The leaders can cooperate by jointly optimizing the cost functionals related to the design stage. Therefore, the cooperative leader problem is as follows:
$\mathscr{P}^{\mathrm{C}}_{\mathrm{leader}}:~{}\min_{\mathbf{a}\in\mathcal{A}}\sum_{j\in\mathcal{M}}J_{j}(a_{j},a_{-j},\{\text{follower strategies}\}).$ (10)
For notation simplicity, let $J(\mathbf{a},\{\text{follower strategies}\})$ be as
$J(\mathbf{a},\cdot)=\sum_{j\in\mathcal{M}}J_{j}(a_{j},a_{-j},\{\text{follower strategies}\}).$
Definition 9 (Cooperative solution for the leaders)
A strategic profile $\tilde{\mathbf{a}}^{*}=(a_{1}^{*},\dots,a_{L}^{*})\in\mathcal{A}$ is a cooperative solution if it solves (10), i.e.,
$\tilde{\mathbf{a}}^{*}\in\arg\min_{\mathbf{a}\in\mathcal{A}}\sum_{j\in\mathcal{M}}J_{j}(a_{j},a_{-j},\{\text{follower strategies}\}).\;\hfill\square$
There are some advantages to solving the problem following a non-cooperative approach for networked systems, e.g., when solving the non-cooperative game problem for the followers, the structure of the controller can be interpreted as a non-centralized controller. Whereas when the followers cooperate, this architecture can be seen as a centralized controller. Therefore, we are interested in judging and measuring how different the non-cooperative and cooperative solutions are. To this end, we use the price of anarchy concept.
II-G Price of Anarchy for Leaders and Followers
The price of anarchy is a key performance indicator that allows measuring how optimal a Nash equilibrium is with respect to the best socially optimal combination [17]. In this work, we take the cost for a Nash equilibrium in comparison to the cost when all the decision-makers cooperate with one another to pursue a social optimum. Indeed, notice that, according to the Definition 2 and Definition 7, there must be multiple Nash equilibrium points. If there are multiple Nash equilibria, then the price of anarchy is computed using the worst equilibrium in terms of its corresponding cost. Suppose you have the set of Nash equilibria for the leaders as $\mathbb{NE}_{\mathrm{leader}}$ and for the followers $\mathbb{NE}_{\mathrm{follower}}$. Then, the corresponding prices of anarchy are defined as follows:
$\mathrm{PoA}_{\mathrm{leader}}$ $=\dfrac{\max\limits_{\mathbf{a}\in\mathbb{NE}_{\mathrm{leader}}}J(\mathbf{a},\mathbf{u})}{\min\limits_{a\in\mathcal{A}}J(\mathbf{a},\mathbf{u})}\geq 1,$ (11a)
$\mathrm{PoA}_{\mathrm{follower}}$ $=\dfrac{\max\limits_{\mathbf{u}\in\mathbb{NE}_{\mathrm{follower}}}V(x_{0},\mathbf{u})}{\min\limits_{u\in\mathcal{U}}V(x_{0},\mathbf{u})}\geq 1.$ (11b)
Note that the denominators in both expressions for the price of anarchy are computed by solving the cooperative game problem presented in Problem (10) and Problem (5). Therefore, the denominators become $J(\tilde{\mathbf{a}}^{*},\{\text{follower strategies}\})$ and $V(x_{0},\tilde{\mathbf{u}}^{*})$, respectively. In addition, let us assume that, as we will see in the numerical example we present in this paper, there is a unique Nash equilibrium for each of the games across the layers, i.e. ${\mathbf{a}}^{*}$ and ${\mathbf{u}}^{*}$. Then, the prices of anarchy can be written as
$\mathrm{PoA}_{\mathrm{leader}}$ $=\dfrac{J(\mathbf{a}^{*},\mathbf{u})}{J(\tilde{\mathbf{a}}^{*},\mathbf{u})},$ $\mathrm{PoA}_{\mathrm{follower}}$ $=\dfrac{V(x_{0},\mathbf{u}^{*})}{V(x_{0},\tilde{\mathbf{u}}^{*})},$
respectively. A lower price of anarchy corresponds to an enhancement in the worst scenario for the equilibrium performance, and $\mathrm{PoA}_{\mathrm{leader}}=1$ or $\mathrm{PoA}_{\mathrm{follower}}=1$ imply that the Nash equilibria are optimal. In other words, when having a unitary price of anarchy, the non-centralized design by the leaders using normal form games, or the non-centralized design of the control law by the followers’ means of dynamic games, are as optimal as those obtained when they jointly optimize in the framework of cooperation.
III Stackelberg equilibrium for
the co-design problem
TABLE I: Different classes of Stackelberg games
Stackelberg Strategic Leaders Interaction Followers Interaction
Game Profile Normal Form Game Difference Game
Class I $(\mathbf{a}^{*},\mathbf{u}^{*})$ Non-cooperative Non-cooperative
Class II $(\tilde{\mathbf{a}}^{*},\tilde{\mathbf{u}}^{*})$ Cooperative Cooperative
Class III $(\mathbf{a}^{*},\tilde{\mathbf{u}}^{*})$ Non-cooperative Cooperative
Class IV $(\tilde{\mathbf{a}}^{*},\mathbf{u}^{*})$ Cooperative Non-cooperative
Figure 1: Different hierarchical strategic interactions illustrating the four Stackelberg game classes (Squares represent leaders and circles represent followers).
As we can consider multiple strategic interactions at each layer of the hierarchical Stackelberg scheme, we have multiple classes for the Stackelberg game as summarized in Table I. Each class is also presented in Fig. 1.
III-A Definition of Stackelberg Equilibrium Classes
This subsection introduces the possible strategic interactions at each layer, which establish a Stackelberg game class. Let us consider the case in which the leaders $\mathcal{L}$ do not cooperate, i.e., there are $L$ designer entities making decisions following an independent interest. Also, assume that once the design is determined, the followers $\mathcal{M}$ design a control action independently, i.e., there are $M$ entities deciding control actions following different non-cooperative costs. The Stackelberg equilibrium corresponding to this scenario is formally presented in Definition 10 below (see Fig. 1).
Definition 10 (Class I: Stackelberg equilibrium with non-cooperative leaders and non-cooperative followers)
A strategic leader-follower profile $(\mathbf{a}^{*},\mathbf{u}^{*})\in\mathcal{A}\times\mathcal{U}$ is a Stackelberg equilibrium for non-cooperative leaders and followers if
$\forall~{}i\in\mathcal{M}:~{}\mathbf{u}_{i}^{*}\in\mathbb{BR}^{\mathcal{M}}_{i}(\mathbf{u}_{-i}^{*},\mathbf{a}^{*}),$ (12a)
$\forall~{}j\in\mathcal{L}:$ (12b)
$a_{j}^{*}\in\arg\min_{a_{j}\in A_{j}}J_{j}\left(a_{j},a_{-j}^{*},\mathbf{u}^{*}\bigg{|}\begin{array}[]{l}\mathbf{u}_{i}^{*}\in\mathbb{BR}^{\mathcal{M}}_{i}(\mathbf{u}_{-i}^{*},a_{j},a_{-j}^{*})\\ \forall i\in\mathcal{M}\end{array}\right).$ (12e)
Note that the leader designs some parameters knowing that the follower is going to optimize based on them. $\square$
Let us now assume that all the leaders $\mathcal{L}$ coordinate each other and agree to jointly perform the design of the system. This cooperation leads to an optimization that is jointly performed by the $L$ leaders. Then, once the system design is established, the followers $\mathcal{M}$ cooperatively design the control actions to operate the system. This cooperative followers’ behavior leads to a centralized control design. Definition 11 formally presents the Stackelberg equilibrium under this scenario (see Fig. 1).
Definition 11 (Class II: Stackelberg equilibrium with cooperative leaders and cooperative followers)
A strategic leader-follower profile $(\tilde{\mathbf{a}}^{*},\tilde{\mathbf{u}}^{*})\in\mathcal{A}\times\mathcal{U}$ is a Stackelberg equilibrium for cooperative leaders and followers if
$\tilde{\mathbf{u}}^{*}$ $\in\mathbb{BR}^{\mathcal{M}}(\tilde{\mathbf{a}}^{*}),$ (13a)
$\tilde{\mathbf{a}}^{*}$ $\in\arg\min_{\mathbf{a}\in\mathcal{A}}\sum_{j\in\mathcal{M}}J_{j}(\mathbf{a},\tilde{\mathbf{u}}^{*}|\tilde{\mathbf{u}}^{*}\in\mathbb{BR}^{\mathcal{M}}(\tilde{\mathbf{a}})).$ (13b)
In this case, all the leaders jointly design parameters against which the followers will optimally respond also jointly by means of cooperation. $\square$
The aforementioned Stackelberg game classes have considered a homogeneous behavior for both the leader and followers, i.e. both layers cooperating or both layers acting independently in a non-cooperative framework. However, layers may exhibit heterogeneous strategic behavior. First, let us consider the case in which the leaders $\mathcal{L}$ do not cooperate for the system design, whereas the followers $\mathcal{M}$ cooperatively react to the system design by solving jointly a centralized optimal control problem. The emerging Stackelberg equilibrium for this scenario is formally presented next in Definition 12 (see Fig. 1).
Definition 12 (Class III: Stackelberg equilibrium with non-cooperative leaders and cooperative followers)
A strategic profile $(\mathbf{a}^{*},\mathbf{u}^{*})\in\mathcal{A}\times\mathcal{U}$ is a Stackelberg equilibrium for non-cooperative leaders and cooperative followers if
$\forall~{}i\in\mathcal{M}:~{}\mathbf{u}_{i}^{*}\in\mathbb{BR}^{\mathcal{M}}_{i}(\mathbf{u}_{-i}^{*},\tilde{\mathbf{a}}^{*}),$
$\tilde{\mathbf{a}}^{*}\in\arg\min_{\mathbf{a}\in\mathcal{A}}\sum_{j\in\mathcal{M}}J_{j}\left(\mathbf{a},\mathbf{u}^{*}|\mathbf{u}_{i}^{*}\in\mathbb{BR}^{\mathcal{M}}_{i}(\mathbf{u}_{-i}^{*},\tilde{\mathbf{a}})~{}\forall i\in\mathcal{M}\right).$
In this case, all the leaders non-cooperatively design the parameters against which the followers will jointly and optimally respond by means of cooperation. $\square$
Finally, let us assume that the leaders $\mathcal{L}$ jointly optimize the system design. Once the system is designed, the followers react to this leader’s strategic action by designing a selfish control input, i.e., the followers $\mathcal{M}$ play a dynamic game. Definition 13 shows he resulting Stackelberg game equilibrium for this combination of behavior across the layers (see Fig. 1).
Definition 13 (Class IV: Stackelberg equilibrium with cooperative leaders and non-cooperative followers)
A strategic profile $(\mathbf{a}^{*},\mathbf{u}^{*})\in\mathcal{A}\times\mathcal{U}$ is a Stackelberg equilibrium for cooperative leaders and non-cooperative followers if
$\tilde{\mathbf{u}}^{*}\in\mathbb{BR}^{\mathcal{M}}(\mathbf{a}^{*}),$
$\forall~{}j\in\mathcal{L}:a_{j}^{*}\in\arg\min_{a_{j}\in A_{j}}J_{j}(a_{j},a_{-j}^{*},\tilde{\mathbf{u}}^{*}|\tilde{\mathbf{u}}^{*}\in\mathbb{BR}^{\mathcal{M}}(a_{j},a_{-j}^{*})).$
In this case, all the leaders jointly design the parameters against which the followers will optimally respond by means of non-cooperative interactions. $\square$
The Stackelberg equilibria for the different classes imply the computation of either a Nash equilibrium given by a fixed point condition or the solution of an optimization problem for the followers in terms of the leaders, and then the same procedure for the leader knowing the optimal response of the followers against the leaders’ decisions. In the following Section, algorithms to solve each one of the classes are presented.
III-B Learning Procedure to Compute Stackelberg Solutions
This section formally presents the procedure to compute each one of the Stackelberg equilibria corresponding to the four different classes introduced above (see Fig. 1). We present a qualitative description of the steps that comprise each one of the computations together with its corresponding algorithm.
III-B1 Class I: Stackelberg game with non-cooperative leaders and non-cooperative followers
Given that under this strategic behavior combination, both sets of decision-makers, i.e., leaders $\mathcal{L}$ and $\mathcal{M}$, play in a non-cooperative manner, the solution for the normal form game and dynamic game problems can be computed by following best-response dynamics. Then, the Stackelberg equilibrium is found for this game class. The steps are the following:
  • For each feasible leader strategic selection $\mathbf{a}\in\mathcal{A}$, let us compute the Nash equilibrium $\mathbf{u}^{*}\in\mathcal{U}$, which is denoted by $\mathbf{u}^{*}(\mathbf{a})$, meaning that the equilibrium strategic profile depends on the strategic selection of the leader. The computed Nash equilibrium is a non-cooperative best response against the leaders’ strategies, i.e., $\mathbf{u}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{NC}}_{\mathcal{M}}(\mathbf{a})$.
  • Using the equilibrium strategic profile $\mathbf{u}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader non-cooperative game problem can be stated only in terms of $\mathbf{a}$, and the Nash equilibrium for the leaders $\mathbf{a}^{*}\in\mathcal{A}$ can be found.
  • Then, using the Nash equilibrium for the leaders game problem $\mathbf{a}^{*}$, the Nash equilibrium for the follower game problem becomes $\mathbf{u}^{*}:=\mathbf{u}^{*}(\mathbf{a}^{*})$.
  • Finally, one obtains the Class I Stackelberg equilibrium as $(\mathbf{u}^{*},\mathbf{a}^{*})\in\mathcal{U}\times\mathcal{A}$.
Details for this computation are presented in Algorithm III-B1.
Algorithm 1 Learning Algorithm for Stackelberg equilibrium in Class I
=$\mathcal{L}=\{1,\dots,L\}$, =$\mathcal{M}=\{1,\dots,M\}$ $A_{j},~{}\forall j\in\mathcal{L}$,
$j\leftarrow 1,~{}i\leftarrow 1,$ $T$
procedure (Non-Cooperative Game for Followers)
     for $k_{1}\leftarrow 1$ to $|A_{1}|$ do
         for $k_{2}\leftarrow 1$ to $|A_{2}|$ do
              $\vdots$
              for $k_{L}\leftarrow 1$ to $|A_{L}|$ do
[…]
                  $\mathbf{a}\leftarrow(a_{1},\dots,a_{L})\in\mathcal{A}$
                  $\mathbf{u}\in\mathcal{U}$
                  while $\mathbf{u}=(\mathbf{u}_{1},\dots,\mathbf{u}_{M})\notin\epsilon-\mathbb{NE}_{\mathrm{follower}}$ do
                       for $i\leftarrow 1$ to $M$ do
                           $\mathbf{u}_{i}\leftarrow\arg\min_{\mathbf{u}_{i}\in\mathcal{U}_{i}}V_{i}(x_{0},\mathbf{u}_{i},\mathbf{u}_{-i})$
                           s. t. (1), $x_{k}\in\mathbb{X}(\mathbf{a}),\forall k\in[0..T]$
                       end for
                  end while
                  $\mathbf{u}^{*}(\mathbf{a})\in\mathcal{U}\leftarrow\mathbf{u}$
              end for
         end for
     end for
end procedure
procedure (Non-Cooperative Game for Leaders)
     while $\mathbf{a}=(a_{1},\dots,a_{L})\notin\epsilon-\mathbb{NE}_{\mathrm{leader}}$ do
         for $j\leftarrow 1$ to $L$ do
              $J_{j}(a_{j},a_{-j})\leftarrow J_{j}(a_{j},a_{-j},\mathbf{u}^{*}(\mathbf{a}))$
              $a_{j}\leftarrow\arg\min_{a_{j}\in A_{j}}J_{j}(a_{j},a_{-j})$
         end for
     end while
     $\mathbf{a}^{*}\leftarrow\mathbf{a}\in\mathcal{A}$
end procedure
procedure (Stackelberg Game Solution)
     $\mathbf{u}^{*}\leftarrow\mathbf{u}^{*}(\mathbf{a}^{*})\in\mathcal{U}$
     $(\mathbf{u}^{*},\mathbf{a}^{*})\in\mathcal{U}\times\mathcal{A}$
end procedure
III-B2 Class II: Stackelberg game with cooperative leaders and cooperative followers
Similar to the previous class, both sets of decision-makers, i.e., leaders $\mathcal{L}$ and $\mathcal{M}$, behave in the same strategic manner. In this combination of strategic actions, both sets across the layers cooperate within their corresponding layer. This structure leads to the computation of a solution for centralized optimization problems. The following are the steps to compute the Stackelberg equilibrium under this game class:
  • For each feasible leader strategic selection $\mathbf{a}\in\mathcal{A}$, let us compute the optimal solution for the cooperative problem given by $\tilde{\mathbf{u}}^{*}\in\mathcal{U}$, which is denoted by $\tilde{\mathbf{u}}^{*}(\mathbf{a})$ as it depends on the strategic selection of the leader. The computed optimal is a cooperative best response against the leaders’ strategies, i.e., $\tilde{\mathbf{u}}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{C}}_{\mathcal{M}}(\mathbf{a})$.
  • Using the optimal strategic profile $\tilde{\mathbf{u}}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader cooperative game problem can be stated only in terms of $\mathbf{a}$, and the optimal solution for the leaders $\tilde{\mathbf{a}}^{*}\in\mathcal{A}$ can be found.
  • Then, using the optimal solution for the leaders game problem $\tilde{\mathbf{a}}^{*}$, the cooperative optimal solution for the follower becomes $\tilde{\mathbf{u}}^{*}:=\tilde{\mathbf{u}}^{*}(\tilde{\mathbf{a}}^{*})$.
  • Finally, one obtains the Class II Stackelberg equilibrium as $(\tilde{\mathbf{u}}^{*},\tilde{\mathbf{a}}^{*})\in\mathcal{U}\times\mathcal{A}$.
Algorithm III-B2 shows the detailed procedure to compute the Stackelberg equilibrium for this class.
Algorithm 2 Learning Algorithm for Stackelberg equilibrium in Class II
=$\mathcal{L}=\{1,\dots,L\}$, =$\mathcal{M}=\{1,\dots,M\}$, $A_{j},~{}\forall j\in\mathcal{L}$
$j\leftarrow 1,~{}i\leftarrow 1,$ $T$
procedure (Control / Cooperative Game for Followers)
     for $k_{1}\leftarrow 1$ to $|A_{1}|$ do
         for $k_{2}\leftarrow 1$ to $|A_{2}|$ do
              $\vdots$
              for $k_{L}\leftarrow 1$ to $|A_{L}|$ do
[…]
                  $\mathbf{a}\leftarrow(a_{1},\dots,a_{L})\in\mathcal{A}$
                  $\tilde{\mathbf{u}}^{*}(\mathbf{a})\leftarrow\arg\min_{\mathbf{u}\in\mathcal{U}}\sum_{i\in\mathcal{M}}V_{i}(x_{0},\mathbf{u})$
                  s. t (1), $x_{k}\in\mathbb{X}(\mathbf{a}),\forall k\in[0..T]$
              end for
         end for
     end for
end procedure
procedure (Control / Cooperative Game for Leaders)
     $J_{j}(\mathbf{a})\leftarrow J_{j}(\mathbf{a},\tilde{\mathbf{u}}^{*}(\mathbf{a}))$
     $\tilde{\mathbf{a}}^{*}\leftarrow\arg\min_{\mathbf{a}\in\mathcal{A}}\sum_{j\in\mathcal{L}}J_{j}(\mathbf{a})$
end procedure
procedure (Stackelberg Game Solution)
     $\tilde{\mathbf{u}}^{*}\leftarrow\tilde{\mathbf{u}}^{*}(\tilde{\mathbf{a}}^{*})\in\mathcal{U}$
     $(\tilde{\mathbf{u}}^{*},\tilde{\mathbf{a}}^{*})\in\mathcal{U}\times\mathcal{A}$
end procedure
III-B3 Class III: Stackelberg game with non-cooperative leaders and cooperative followers
This class comprises a heterogeneous strategic behavior for the leaders $\mathcal{L}$ with respect to the followers $\mathcal{M}$. In this class, the leaders behave selfishly as they independently optimize a cost function, i.e., this layer is solved by following a best-response dynamics to compute the corresponding solution in the normal form game. Then, as the followers jointly optimize to compute the control input, this step consists of a standard centralized optimal control problem. Next, we present the steps to compute the Stackelberg equilibrium for this class.
  • For each feasible leader strategic selection $\mathbf{a}\in\mathcal{A}$, let us compute the optimal solution for the cooperative problem that corresponds to a control problem given by $\tilde{\mathbf{u}}^{*}\in\mathcal{U}$, which can be denoted by $\tilde{\mathbf{u}}^{*}(\mathbf{a})$ as it depends on the strategic selection of the leader. The computed optimal is a cooperative best response against the leaders’ strategies, i.e., $\tilde{\mathbf{u}}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{C}}_{\mathcal{M}}(\mathbf{a})$.
  • Using the optimal strategic profile $\tilde{\mathbf{u}}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader cooperative game problem can be stated only in terms of $\mathbf{a}$, and the Nash equilibrium for the leaders $\mathbf{a}^{*}\in\mathcal{A}$ can be found.
  • Then, using the Nash equilibrium solution for the leaders game problem $\mathbf{a}^{*}$, the cooperative optimal solution for the follower becomes $\tilde{\mathbf{u}}^{*}:=\tilde{\mathbf{u}}^{*}(\mathbf{a}^{*})$.
  • Finally, one obtains the Class III Stackelberg equilibrium as $(\tilde{\mathbf{u}}^{*},\mathbf{a}^{*})\in\mathcal{U}\times\mathcal{A}$.
Algorithm III-B3 shows the procedure to compute the Stackelberg equilibrium for this class, comprising a game problem for the leaders and a centralized control problem for the followers.
Algorithm 3 Learning Algorithm for Stackelberg equilibrium in Class III
=$\mathcal{L}=\{1,\dots,L\}$, =$\mathcal{M}=\{1,\dots,M\}$, $A_{j},~{}\forall j\in\mathcal{L}$
$j\leftarrow 1,~{}i\leftarrow 1,$ $T$
procedure (Control / Cooperative Game for Followers)
     for $k_{1}\leftarrow 1$ to $|A_{1}|$ do
         for $k_{2}\leftarrow 1$ to $|A_{2}|$ do
              $\vdots$
              for $k_{L}\leftarrow 1$ to $|A_{L}|$ do
[…]
                  $\mathbf{a}\leftarrow(a_{1},\dots,a_{L})\in\mathcal{A}$
                  $\tilde{\mathbf{u}}^{*}(\mathbf{a})\leftarrow\arg\min_{\mathbf{u}\in\mathcal{U}}\sum_{i\in\mathcal{M}}V_{i}(x_{0},\mathbf{u})$
                  s. t. (1), $x_{k}\in\mathbb{X}(\mathbf{a}),\forall k\in[0..T]$
              end for
         end for
     end for
end procedure
procedure (Non-Cooperative Game for Leaders)
     while $\mathbf{a}=(a_{1},\dots,a_{L})\notin\epsilon-\mathbb{NE}_{\mathrm{leader}}$ do
         for $j\leftarrow 1$ to $L$ do
              $J_{j}(a_{j},a_{-j})\leftarrow J_{j}(a_{j},a_{-j},\tilde{\mathbf{u}}^{*}(\mathbf{a}))$
              $a_{j}\leftarrow\arg\min_{a_{j}\in A_{j}}J_{j}(a_{j},a_{-j})$
         end for
     end while
     $\mathbf{a}^{*}\leftarrow\mathbf{a}\in\mathcal{A}$
end procedure
procedure (Stackelberg Game Solution)
     $\tilde{\mathbf{u}}^{*}\leftarrow\mathbf{u}^{*}(\mathbf{a}^{*})\in\mathcal{U}$
     $(\tilde{\mathbf{u}}^{*},\mathbf{a}^{*})\in\mathcal{U}\times\mathcal{A}$
end procedure
III-B4 Class IV: Stackelberg game with cooperative leaders and non-cooperative followers
The last possible combination of strategic behaviors consists of allowing the leaders to cooperate, meaning that this layer game problem is solved by means of solving an optimization problem. Then, given that the followers $\mathcal{M}$ do not cooperate, this layer is solved by means of a best-response dynamics in the context of dynamic games. The following shows the required steps for the computation of the Stackelberg equilibrium:
  • For each feasible leader strategic selection $\mathbf{a}\in\mathcal{A}$, let us compute the Nash equilibrium $\mathbf{u}^{*}\in\mathcal{U}$, which is denoted by $\mathbf{u}^{*}(\mathbf{a})$, meaning that the equilibrium strategic profile depends on the strategic selection of the leader. The computed Nash equilibrium is a non-cooperative best response against the leaders’ strategies, i.e., $\mathbf{u}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{NC}}_{\mathcal{M}}(\mathbf{a})$.
  • Using the equilibrium strategic profile $\mathbf{u}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader cooperative problem can be stated only in terms of $\mathbf{a}$, and the joint optimization for the leaders can be computed, i.e., $\tilde{\mathbf{a}}^{*}\in\mathcal{A}$ can be found.
  • Then, using the cooperative solution for the leaders $\tilde{\mathbf{a}}^{*}$, the Nash equilibrium for the follower game problem becomes $\mathbf{u}^{*}:=\mathbf{u}^{*}(\tilde{\mathbf{a}}^{*})$.
  • Finally, one obtains the Class IV Stackelberg equilibrium as $(\mathbf{u}^{*},\tilde{\mathbf{a}}^{*})\in\mathcal{U}\times\mathcal{A}$.
Algorithm III-B4 shows the detailed procedure to compute the Stackelberg equilibrium for this class, comprising a dynamic game problem for the followers and an optimization problem for the leaders.
Algorithm 4 Learning Algorithm for Stackelberg equilibrium in Class IV
=$\mathcal{L}=\{1,\dots,L\}$, =$\mathcal{M}=\{1,\dots,M\}$, $A_{j},~{}\forall j\in\mathcal{L}$
$j\leftarrow 1,~{}i\leftarrow 1,$ $T$
procedure (Non-Cooperative Game for Followers)
     for $k_{1}\leftarrow 1$ to $|A_{1}|$ do
         for $k_{2}\leftarrow 1$ to $|A_{2}|$ do
              $\vdots$
              for $k_{L}\leftarrow 1$ to $|A_{L}|$ do
[…]
                  $\mathbf{a}\leftarrow(a_{1},\dots,a_{L})\in\mathcal{A}$
                  $\mathbf{u}\in\mathcal{U}$
                  while $\mathbf{u}=(\mathbf{u}_{1},\dots,\mathbf{u}_{M})\notin\epsilon-\mathbb{NE}_{\mathrm{follower}}$ do
                       for $i\leftarrow 1$ to $M$ do
                           $\mathbf{u}_{i}\leftarrow\arg\min_{\mathbf{u}_{i}\in\mathcal{U}_{i}}V_{i}(x_{0},\mathbf{u}_{i},\mathbf{u}_{-i})$
                           s. t.  (1), $x_{k}\in\mathbb{X}(\mathbf{a}),\forall k\in[0..T]$
                       end for
                  end while
                  $\mathbf{u}^{*}(\mathbf{a})\in\mathcal{U}\leftarrow\mathbf{u}$
              end for
         end for
     end for
end procedure
procedure (Control / Cooperative Game for Leaders)
     $J_{j}(\mathbf{a})\leftarrow J_{j}(\mathbf{a},\mathbf{u}^{*}(\mathbf{a}))$
     $\tilde{\mathbf{a}}^{*}\leftarrow\arg\min_{\mathbf{a}\in\mathcal{A}}\sum_{j\in\mathcal{L}}J_{j}(\mathbf{a})$
end procedure
procedure (Stackelberg Game Solution)
     $\mathbf{u}^{*}\leftarrow\mathbf{u}^{*}(\tilde{\mathbf{a}}^{*})\in\mathcal{U}$
     $(\mathbf{u}^{*},\tilde{\mathbf{a}}^{*})\in\mathcal{U}\times\mathcal{A}$
end procedure
After having introduced the theoretical foundation of the bi-level game problems for leaders and followers interacting in the framework of Stackelberg games, we present, in the following section, a networked system to illustrate an engineering application for the co-design problem.
IV Networked System Application
The Barcelona Drinking Water Network (DWN) in [18, Chapter 3, Fig. 3.4], managed by the company Aguas de Barcelona (AGBAR), supplies drinking water to Barcelona city and its metropolitan area, utilizing water from the Ter and Llobregat rivers—regulated upstream by dams with a combined reservoir capacity of 600 hm³, alongside groundwater from the Besós River aquifer and supplementary wells. The network integrates four drinking water treatment plants: the Abrera and Sant Joan Despì facilities (Llobregat River), the Cardedeu plant (Ter River), and the Besós plant (groundwater), with additional pumping infrastructure extracting from wells to achieve a total flow of approximately 7 m³/s. This is a well-known benchmark that has been used to illustrate networked control applications, game-theory-based control analysis, and optimization-based controllers. Here, we use this case study to evaluate and illustrate the Stackelberg game classes introduced in this paper.
Let us consider the Barcelona drinking water network as a networked multi-agent system to be studied for the Stackelberg game for the co-design problem. Let $f({x}_{k},{u}_{1,k},\dots,{u}_{M,k})=f({x}_{k},\{{u}_{i,k}\}_{i\in\mathcal{M}})$ be a linear system as follows:
${x}_{k+1}$ $=A{x}_{k}+\sum_{i\in\mathcal{M}}B_{i}{u}_{i,k}+B_{l}{d}_{k},~{}\forall k\in[0..(T-1)],$ (14a)
$x_{0}$ $\in\mathbb{X}(\cdot)~{}\text{given},$ (14b)
where ${x}\in\mathbb{R}^{17}$ denotes the system states corresponding to the water level at each one of the reservoirs in the DWN, ${u}\in\mathbb{R}^{61}$ denotes the control inputs corresponding to the controllable flows determined by either valves or pumps throughout the DWN, and ${d}\in\mathbb{R}^{25}$ denotes the vector of time-varying demands, which are assumed to be obtained using forecasting methodologies. The feasible set for the system states is
$\mathbb{X}(\cdot)$ =$=\{{x}\in\mathbb{R}^{17}:{x}^{\mathrm{min}}\leq{x}\leq{x}^{\mathrm{max}}(\cdot)\},$
where $\mathbb{X}(\cdot):=\mathbb{X}(\{\text{designed parameter}\})$ , ${x}^{\mathrm{max}}(\cdot):={x}^{\mathrm{max}}(\{\text{designed parameter}\})$. $\bar{{x}}^{\mathrm{max}}\in\mathbb{R}^{17}$ be a nominal parameter for the system state constraints, i.e., when there is no co-design problem under consideration, then ${x}^{\mathrm{max}}=\bar{{x}}^{\mathrm{max}}$.
In addition, the feasible set for the control inputs is
$\mathbb{U}=\{{u}\in\mathbb{R}^{61}:{u}^{\mathrm{min}}\leq{u}\leq{u}^{\mathrm{max}}\}.$
IV-A Leaders and Followers in the Barcelona DWN
In terms of the co-design problem of the Barcelona DWN, we consider $L=4$ leaders within the system in charge of the design of four reservoirs’ dimensions, i.e., =$\mathcal{L}=\{1,\dots,4\}$ corresponding to the states $x_{1},x_{4},x_{10},x_{14}$, respectively (see Table II). Each leader in $\mathcal{L}$ has a set of possible actions in the strategic game
$A_{j}=\{0.5,0.75,1,1.25,1.5\},~{}\forall~{}j\in\mathcal{L}.$
Also, we have that $x_{1}^{\mathrm{max}}=a_{1}\cdot\bar{x}_{1}^{\mathrm{max}}$, $x_{4}^{\mathrm{max}}=a_{2}\cdot\bar{x}_{4}^{\mathrm{max}}$, $x_{10}^{\mathrm{max}}=a_{3}\cdot\bar{x}_{10}^{\mathrm{max}}$, $x_{14}^{\mathrm{max}}=a_{4}\cdot\bar{x}_{14}^{\mathrm{max}}$, where $\bar{x}_{1}^{\mathrm{max}}$, $\bar{x}_{4}^{\mathrm{max}}$, $\bar{x}_{10}^{\mathrm{max}}$, and $\bar{x}_{14}^{\mathrm{max}}$ are the nominal values for the reservoirs under design. For all the other system state maximum values, we have that $x_{j}^{\mathrm{max}}=\bar{x}_{j}^{\mathrm{max}}$ with $j\in\{1,\dots,17\}\setminus\{1,4,10,14\}$.
TABLE II: Decision-makers $\{1,2,3,4\}$ corresponding to sub-systems.
No. of Subsystem Color Total Amount of States Designed State
1 Green 4 $x_{4}$
2 Blue 8 $x_{10}$
3 Magenta 3 $x_{14}$
4 Red 2 $x_{1}$
On the other hand, let us consider $M=4$ followers, i.e., =$\mathcal{M}=\{1,\dots,4\}$, which are in charge of four different sub-systems of the DWN. As shown in Fig. 2, four subsystems highlighted by different colors are considered. This partitioning has been adopted from the research reported in [19]. Thus, each follower decides the control inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{4}$ corresponding to the sub-systems with colors green, blue, magenta, and red, respectively.
Note that the strategic selection of the followers $\mathcal{M}$ is influenced by the strategic selection of the leaders $\mathcal{L}$. When the leaders decide on a design for the reservoirs at each one of the sub-systems, the followers modify accordingly their optimal strategies. In the following, we introduce the game theoretical problems for both the leaders and followers, considering either cooperative or non-cooperative behavior. Let us start by introducing the normal form game problems for the leaders.
Refer to caption Refer to caption
(c) (d)
Figure 2: Decision-makers in the followers’ layer. (a) Decision-maker 1, (b) Decision-maker 2, (c) Decision-maker 3, (d) Decision-maker 4.
IV-B Non-Cooperative and Cooperative Leaders Game
Let the leaders in $\mathcal{L}$ behave in a non-cooperative manner, i.e., each leader decides its strategies independently. The non-cooperative game problem is given by
$\forall~{}j\in\mathcal{L}$ $:~{}~{}\min_{a_{j}\in A_{j}}J_{j}(a_{j},a_{-j},\mathbf{u}),$
with a leader cost functional of the form:
$J_{j}(a_{j},a_{-j},\mathbf{u})$ $=g(\mathbf{a})+h(\mathbf{u}),~{}\forall~{}j\in\mathcal{L},$
where $g:\mathcal{A}\to\mathbb{R}$ and $h:\mathcal{U}\to\mathbb{R}$. For example, we may consider the following functions for the leader costs:
$g(\mathbf{a})$ $=\mathbf{a}^{\top}Q_{j}\mathbf{a}+v_{j}^{\top}\mathbf{a},$
$h(\mathbf{u})$ $=\sum_{i\in\mathcal{M}}\sum_{k=0}^{T}\alpha_{i,k}^{\top}{u}_{i,k},$
where $Q_{j}\succeq 0$ and $S_{ij}\succeq 0$ are weighting matrices, $v_{j}$ is a given vector. For leader costs, the parameters are chosen as $Q_{j}=0$, $v_{j}=0.01$, and $\alpha_{i,k}$ is the vector of time-varying electricity prices per input unit for follower $i$ at time $k$. The function $g(\mathbf{a})$ is used to penalize the effort that the leader applies in the design of the reservoirs. This cost can be associated with economic costs for implementing the design. In this regard, the leader is interested in minimizing the magnitude of its strategic selection. On the other hand, the cost function $h(\mathbf{u}$ depends on the followers’ strategic selection. This means that, when the leader makes decisions over the modeling, it also takes into consideration how the followers will perform their control actions. Note that this is a game problem as the cost functional of the $j-$th decision-maker (leader) is affected by the decisions made by $\mathcal{L}\setminus\{j\}$ through the followers’ actions $\mathbf{u}_{i}$, for all $i\in\mathcal{M}$. To emphasize this coupling, notice that the strategic design of a single leader, e.g., the $j-$th leader, affects the evolution of the control actions for all the followers $\mathcal{M}$ as they are dynamically coupled through (1). Therefore, as all the leaders $\mathcal{L}$ take into consideration all the followers’ actions in their cost functionals, the decisions of the $j-$th leader affects the cost of all the other leaders $\mathcal{L}\setminus\{j\}$ through $\mathbf{u}_{i}$ for all $i\in\mathcal{M}$.
Now, let us assume that all the leaders $\mathcal{L}$ agree on cooperating in the design of the system. Therefore, all the leaders jointly solve the following optimization problem:
$\min_{\mathbf{a}\in\mathcal{A}}\sum_{j\in\mathcal{L}}J_{j}(a_{j},a_{-j},\mathbf{u}).$ (15)
This problem can be interpreted as a direct optimization of the design parameters, taking into consideration how the followers will react against the design.
IV-C Non-Cooperative and Cooperative Followers Game
We next introduce the specific dynamic game problem for the followers in the DWN. Each follower decision-maker deciding over each sub-system of the DWN performs in a non-cooperative fashion. Then, the dynamic game problem for the DWN is as follows:
$\forall i\in\mathcal{M}:\min_{\mathbf{u}_{i}\in\mathcal{U}_{i}}V_{i}(x_{0},\mathbf{u}_{i},\mathbf{u}_{-i}),$
[…]
where the cost for each follower is given by
$V_{i}(x_{0},\mathbf{u})=\sum_{k=0}^{T}\alpha_{i,k}^{\top}{u}_{i,k}+\Delta{u}_{i,k}^{\top}R_{i}\Delta{u}_{i,k},~{}\forall i\in\mathcal{M},$ (17)
where $\Delta{u}_{i,k}={u}_{i,k}-{u}_{i,k-1}$ and $R_{i}\succ 0$, and $T>0$ is a planning horiizon. The feasible sets for both the control strategies and system states are as follows:
$\mathbb{U}_{i}$ :=$:=\{{u}_{i}\in\mathbb{R}^{n_{u_{i}}}:{u}^{\mathrm{min}}_{i}\leq{u}_{i}\leq{u}^{\mathrm{max}}_{i}\},$ (18)
$\mathbb{X}(\mathbf{a})$ :=$:=\{{x}\in\mathbb{R}^{n_{x}}:{x}^{\mathrm{min}}\leq{x}\leq{x}^{\mathrm{max}}(\mathbf{a})\}.$ (19)
It is important to highlight that the leaders’ decisions directly affect the system state constraints for each one of the followers. In addition, note that modifying a single reservoir’s constraint has an impact over the whole networked system, i.e., over the whole set of decision-makers $\mathcal{M}$ given the constraint given by the dynamical system in (1).
(a) Subsystem 1
(b) Subsystem 2
(c) Subsystem 3
(d) Subsystem 4
Figure 3: Optimal follower solutions with four classes.
(a) Subsystem 1
(b) Subsystem 2
(c) Subsystem 3
(d) Subsystem 4
Figure 4: State evaluations with optimal leader and follower solutions.
If the followers decide to cooperate in order to define the appropriate control inputs, then the problem becomes a traditional optimal control problem. The cooperative game problem is as follows:
$\min_{(\mathbf{u}_{1},\dots,\mathbf{u}_{M})\in\prod_{i\in\mathcal{M}}\mathcal{U}_{i}}\sum_{i\in\mathcal{M}}V_{i}(x_{0},\mathbf{u}),$
[…]
where $B=[B_{1}\quad\dots\quad B_{M}]$, $E=[E_{1}\quad\dots\quad E_{M}]$, and ${u}_{k}=[{u}_{1,k}^{\top}\quad\dots\quad{u}_{M,k}^{\top}]^{\top}$. We compute and test each one of the Stackelberg game classes presented in Fig. 1 by combining the aforementioned game problems. The results are presented and discussed in the coming section, where we present the Stackelberg equilibrium for each class and we also analyze the price of anarchy at each layer (leader and follower layer).
V Results and Discussions
TABLE III: Optimal leader solutions with four classes.
Leader Follower $\mathbf{a}^{*}$ $J_{1}^{*}$ $J_{2}^{*}$ $J_{3}^{*}$ $J_{4}^{*}$ $J^{*}$
Class I Non-cooperative Non-cooperative [0.5, 0.5, 0.5, 1]0.50.50.51[0.5,\;0.5,\;0.5,\;1][ 0.5 , 0.5 , 0.5 , 1 ] 252,647 252,711 252,819 252,635 1,010,812
Class II Cooperative Cooperative [1.5, 1.25, 0.5, 1.5]1.51.250.51.5[1.5,\;1.25,\;0.5,\;1.5][ 1.5 , 1.25 , 0.5 , 1.5 ] 250,340 250,491 250,479 250,297 1,001,607
Class III Non-cooperative Cooperative [0.5, 0.5, 0.5, 0.5]$0.5$[0.5,\;0.5,\;0.5,\;0.5][ 0.5 , 0.5 , 0.5 , 0.5 ] 250,386 250,449 250,558 250,372 1,001,765
Class IV Cooperative Non-cooperative [1, 0.75, 0.5, 1]10.750.51[1,\;0.75,\;0.5,\;1][ 1 , 0.75 , 0.5 , 1 ] 252,620 252,707 252,776 252,592 1,010,695
TABLE IV: Optimal follower solutions with four classes.
Leader Follower $\mathbf{a}^{*}$ $V_{1}^{*}$ $V_{2}^{*}$ $V_{3}^{*}$ $V_{4}^{*}$ $V^{*}$
Class I Non-cooperative Non-cooperative [0.5, 0.5, 0.5, 1]0.50.50.51[0.5,\;0.5,\;0.5,\;1][ 0.5 , 0.5 , 0.5 , 1 ] 0.0732 17.4186 18.3429 0.7798 36.6145
Class II Cooperative Cooperative [1.5, 1.25, 0.5, 1.5]1.51.250.51.5[1.5,\;1.25,\;0.5,\;1.5][ 1.5 , 1.25 , 0.5 , 1.5 ] 0.0743 17.0877 17.5127 1.7343 36.4090
Class III Non-cooperative Cooperative [0.5, 0.5, 0.5, 0.5]$0.5$[0.5,\;0.5,\;0.5,\;0.5][ 0.5 , 0.5 , 0.5 , 0.5 ] 0.0740 17.0852 17.5375 1.7099 36.4066
Class IV Cooperative Non-cooperative [1, 0.75, 0.5, 1]10.750.51[1,\;0.75,\;0.5,\;1][ 1 , 0.75 , 0.5 , 1 ] 0.0716 17.3884 18.3939 0.7383 36.5923
TABLE V: Price of Anarchy computation results.
Scenario $\mathrm{PoA}$
Leader Cooperative followers 1.0002
Non-cooperative followers 1.0001
Follower Cooperative leaders 1.0050
Non-cooperative leaders 1.0057
All simulations were conducted over a 72-hour period (3 days) with a sampling time interval of 1 hour. Water demands and electricity prices exhibit distinct daily patterns. The selected optimal solutions of the followers are shown in Fig. 3. In Figs. 3(a) and 3(b), the optimal follower solutions vary significantly, primarily attributable to differences in tank designs across four distinct classes and the implementation of two different control strategies (cooperative versus non-cooperative). Specifically, Fig. 3(b) highlights the scenarios where the inputs occasionally reach the maximum values because identical input constraints are applied. In Fig. 3(c), due to the same leader solutions with four classes, the same optimal follower solutions are observed. Fig. 3(d) further shows the impact of two control strategies: the follower solutions with Class I and Class IV, both governed by a non-cooperative control framework, exhibit striking similarities, as do those for Classes II and III, which adopt a cooperative control approach. Moreover, it can also be observed that the follower solutions, i.e., the flows through actuators (valves and pumps), have a potential daily pattern due to the water demand satisfaction.
Fig. 4 illustrates the volume evolutions of selected tanks across four subsystems, each corresponding to one of the four classes. The observed daily pattern in tank volumes mirrors the water demand cycle. In Fig. 4(c), the volume evolutions of tank $x_{14}$ are similar to the four classes, as the designed tank sizes are the same based on the leader solutions, which is consistent with the follower solutions shown in Fig. 3(c). From the game theory perspective, the trajectories presented in Fig. 3(a) and Fig. 3(d) correspond to the Nash equilibrium for the followers when they strategically interact in a dynamic game (Classes I and IV). In contrast, the trajectories in Fig. 3(b) and Fig. 3(c) present the optimal control inputs corresponding to a cooperative dynamic game (Classes II and III). The reader may compare the followers’ strategic interactions in Fig. 1.
[…]
The optimal costs for both leaders and followers, and for all the possible classes (from Class I to Class IV), are presented in Table III and Table IV. By using such optimal values, one can measure or evaluate the difference between cooperating and non-cooperating, or from the control perspective, one can evaluate the cost difference of a centralized controller in front of a decentralized controller. We perform this assessment by means of the price-of-anarchy introduced in Section II. The results of the price-of-anarchy corresponding to all the interactive combinations are presented in Table V. It is interesting to observe that the price-of-anarchy is quite close to one in all the cases, indicating that the obtained Nash equilibria are optimal. In other words, the decentralized control approach design is optimal.
VI Concluding Remarks and Future Directions
We have presented multiple classes of Stackelberg games for the co-design of networked systems comprising the simultaneous design of both system and control. Under this approach, a leader is in charge of deciding on a design system parameter. Note that, in general, this decision can be related to the selection of elements such as actuators, or any other system specification. Then, there is a follower layer where the control design takes place. As shown in this paper, the decisions made at the control design depend on the decisions made at the system design stage. Hence, we have presented multiple possibilities for such a bi-level Stackelberg-like interaction. This is because we can consider multiple parties at each one of the layers, i.e., multiple leaders and multiple followers, leading to more involved game-theoretical settings. We consider the case in which leaders and followers can either cooperate or not, and all the possible combinations for these interactions. We have shown that the cooperative game approach coincides with a control problem, and the non-cooperative game can be seen as a decentralized control strategy. Moreover, the evaluation of the price-of-anarchy for the computed equilibrium solutions shows that the decentralized controllers are optimal. Finally, as it was highlighted in the manuscript, we have considered a discrete finite set of strategies for the leaders. As future work, it is proposed to consider a continuum set for the strategic leader set.
Acknowledgments
Julian Barreiro-Gomez is profoundly grateful to God and Our Lady of Lourdes for the blessings of health and life, without whose intercession this work would not have been possible. Ye Wang acknowledges support from the Australian Research Council through the Discovery Early Career Researcher Award (DE220100609).